This file, cyinpenc.dtx, is to be included into inputenc.dtx in a future L^{A}TEX release. It is part of the T2 package.

1 The Cyrillic codepages

There are several widely used Russian (and, generally, Cyrillic) codepages. Currently, we define here the following codepages:

• cp 866 is the standard MS-DOS Russian codepage. There are also several codepages in use, which are very similar to the cp 866. These are: so called “Cyrillic Alternative codepage” (or Alternative Variant of cp 866), Modified Alternative Variant, and the New Alternative Variant. The differences take place in the range 0xf2–0xfe. These ‘Alternative’ codepages are also supported.

• cp 855 is the standard MS-DOS Cyrillic codepage.

• cp 1251 is the standard MS Windows Cyrillic codepage.

• koi8-r is a standard codepage widely used in UNIX-like systems for Russian language support. It is specified in RFC 1489. The situation with koi8-r is somewhat similar to the one with cp 866: there are also several similar codepages in use, which coincide with koi8-r for all Russian letters, but add some other cyrillic letters. These codepages include: koi8-u (it is a variant of the koi8-r codepage with some Ukrainian letters added), koi8-ru (it is described in a draft RFC document specifying the widely used character set for mail and news exchange in the Ukrainian internet community as well as for presentation WWW information resources in Ukrainian language), and ISO-IR-111 ECMA Cyrillic Code Page. All these codepages are also supported.

• ISO 8859-5 Cyrillic codepage (also ISO-IR-144).

• Apple Macintosh Cyrillic (Microsoft cp 10007) codepage.

• Apple Macintosh Ukrainian codepage (very similar to the previous codepage).

For all codepages, X2 (or T2*) encoding is needed. For cp 1251 and both Apple Macintosh encodings, T1 is necessary also. However, if the characters used from these codepages will be limited only to Russian letters, it is sufficient to have old LH fonts with LCY or LWN encoding. In this case the characters which are absent in the font will cause error messages.

Note, that the following glyphs constructed using accents are not ‘named’ here:

\CYRGJE (\'\CYRG), \cyrgje (\'\cyrg), \CYRKJE (\'\CYRK), \cyrkje (\'\cyrk).

Also, \@tabacckludge' is used instead of \' because of tabbing environment.

1.1 Headers

 1 ⟨cp866⟩\ProvidesFile{cp866.def}
 2 ⟨cp866av⟩\ProvidesFile{cp866av.def}
 3 ⟨cp866mav⟩\ProvidesFile{cp866mav.def}
 4 ⟨cp866nav⟩\ProvidesFile{cp866nav.def}
 5 ⟨cp1251⟩\ProvidesFile{cp1251.def}
 6 ⟨cp855⟩\ProvidesFile{cp855.def}
 7 ⟨koi8 & koi8r⟩\ProvidesFile{koi8-r.def}
 8 ⟨koi8 & koi8ru⟩\ProvidesFile{koi8-ru.def}
 9 ⟨koi8 & isoir111⟩\ProvidesFile{isoir111.def}
10 ⟨koi8 & koi8u⟩\ProvidesFile{koi8-u.def}
11 ⟨ISO88595⟩\ProvidesFile{8859-5.def}
12 ⟨maccyrillic⟩\ProvidesFile{maccyr.def}
13 ⟨macukrainian⟩\ProvidesFile{macukr.def}
14 [1998/05/12 v1.0 Input encoding file]
15 %\NeedsTeXFormat{LaTeX2e}[1995/12/01]

16 ⟨∗cp866 | cp855 | koi8r | koi8u⟩

17 \ProvideTextCommandDefault{\textblacksquare}
18 {\vrule \@width .3em \@height .4em \@depth -.1em\relax}

19 ⟨/cp866 | cp855 | koi8r | koi8u⟩

20 ⟨cp866 | cp1251⟩\providecommand{\mathdegree}{{^\circ}}

21 ⟨∗cp1251 | koi8ru | maccyrillic | macukrainian⟩

22 \ProvideTextCommandDefault{\texttrademark}
23 {\UseTextSymbol{T1}{\texttrademark}}
24 \ProvideTextCommandDefault{\textregistered}
25 {\UseTextSymbol{T1}{\textcircled{\scshape R}}}
26 \ProvideTextCommandDefault{\textcopyright}
27 {\UseTextSymbol{T1}{\textcircled{c}}}

28 ⟨/cp1251 | koi8ru | maccyrillic | macukrainian⟩ 29 ⟨∗cp1251⟩

30 \ProvideTextCommandDefault{\quotesinglbase}
31 {\UseTextSymbol{T1}{\quotesinglbase}}
32 \ProvideTextCommandDefault{\textbrokenbar}
33 {\TextSymbolUnavailable\textbrokenbar}
34 \ProvideTextCommandDefault{\texteuro}
35 {\TextSymbolUnavailable\texteuro}

36 ⟨/cp1251⟩

37 ⟨koi8r | koi8u⟩\providecommand{\mathtwosuperior}{{^2}}
38 ⟨cp866 & MAV⟩\providecommand{\mathnsuperior}{{^n}}

1.2 Microsoft cp 866

39 ⟨∗cp866⟩

40 \DeclareInputText{128}{\CYRA}
41 \DeclareInputText{129}{\CYRB}
42 \DeclareInputText{130}{\CYRV}
43 \DeclareInputText{131}{\CYRG}
44 \DeclareInputText{132}{\CYRD}
45 \DeclareInputText{133}{\CYRE}
46 \DeclareInputText{134}{\CYRZH}

47 \DeclareInputText{135}{\CYRZ}
48 \DeclareInputText{136}{\CYRI}
49 \DeclareInputText{137}{\CYRISHRT}
50 \DeclareInputText{138}{\CYRK}
51 \DeclareInputText{139}{\CYRL}
52 \DeclareInputText{140}{\CYRM}
53 \DeclareInputText{141}{\CYRN}
54 \DeclareInputText{142}{\CYRO}
55 \DeclareInputText{143}{\CYRP}
56 \DeclareInputText{144}{\CYRR}
57 \DeclareInputText{145}{\CYRS}
58 \DeclareInputText{146}{\CYRT}
59 \DeclareInputText{147}{\CYRU}
60 \DeclareInputText{148}{\CYRF}
61 \DeclareInputText{149}{\CYRH}
62 \DeclareInputText{150}{\CYRC}
63 \DeclareInputText{151}{\CYRCH}
64 \DeclareInputText{152}{\CYRSH}
65 \DeclareInputText{153}{\CYRSHCH}
66 \DeclareInputText{154}{\CYRHRDSN}
67 \DeclareInputText{155}{\CYRERY}
68 \DeclareInputText{156}{\CYRSFTSN}
69 \DeclareInputText{157}{\CYREREV}
70 \DeclareInputText{158}{\CYRYU}
71 \DeclareInputText{159}{\CYRYA}

72 %

73 \DeclareInputText{160}{\cyra}
74 \DeclareInputText{161}{\cyrb}
75 \DeclareInputText{162}{\cyrv}
76 \DeclareInputText{163}{\cyrg}
77 \DeclareInputText{164}{\cyrd}
78 \DeclareInputText{165}{\cyre}
79 \DeclareInputText{166}{\cyrzh}
80 \DeclareInputText{167}{\cyrz}
81 \DeclareInputText{168}{\cyri}
82 \DeclareInputText{169}{\cyrishrt}
83 \DeclareInputText{170}{\cyrk}
84 \DeclareInputText{171}{\cyrl}
85 \DeclareInputText{172}{\cyrm}
86 \DeclareInputText{173}{\cyrn}
87 \DeclareInputText{174}{\cyro}
88 \DeclareInputText{175}{\cyrp}
89 \DeclareInputText{224}{\cyrr}
90 \DeclareInputText{225}{\cyrs}
91 \DeclareInputText{226}{\cyrt}
92 \DeclareInputText{227}{\cyru}
93 \DeclareInputText{228}{\cyrf}
94 \DeclareInputText{229}{\cyrh}
95 \DeclareInputText{230}{\cyrc}
96 \DeclareInputText{231}{\cyrch}

 97 \DeclareInputText{232}{\cyrsh}
 98 \DeclareInputText{233}{\cyrshch}
 99 \DeclareInputText{234}{\cyrhrdsn}
100 \DeclareInputText{235}{\cyrery}
101 \DeclareInputText{236}{\cyrsftsn}
102 \DeclareInputText{237}{\cyrerev}
103 \DeclareInputText{238}{\cyryu}
104 \DeclareInputText{239}{\cyrya}

105 %

106 \DeclareInputText{240}{\CYRYO}
107 \DeclareInputText{241}{\cyryo}

The following block corresponds to the standard cp 866 codepage:

108 ⟨∗std⟩

109 \DeclareInputText{242}{\CYRIE}
110 \DeclareInputText{243}{\cyrie}
111 \DeclareInputText{244}{\CYRYI}
112 \DeclareInputText{245}{\cyryi}
113 \DeclareInputText{246}{\CYRUSHRT}
114 \DeclareInputText{247}{\cyrushrt}

115 %

116 \DeclareInputMath{248}{\mathdegree}
117 \DeclareInputText{249}{\textbullet}
118 \DeclareInputText{250}{\textperiodcentered}
119 \DeclareInputMath{251}{\surd}
120 \DeclareInputText{252}{\CYRNo}
121 \DeclareInputText{253}{\textcurrency}
122 \DeclareInputText{254}{\textblacksquare}

123 ⟨/std⟩

The following block corresponds to the so called ‘Alternative Variant’ (AV) of cp 866:

124 ⟨∗AV⟩

125 % 0xf2 LOW ACUTE ACCENT
126 % 0xf3 LOW GRAVE ACCENT
127 % 0xf4 HIGH ACUTE ACCENT
128 % 0xf5 HIGH GRAVE ACCENT
129 \DeclareInputMath{246}{\rightarrow}
130 \DeclareInputMath{247}{\leftarrow}
131 \DeclareInputMath{248}{\downarrow}
132 \DeclareInputMath{249}{\uparrow}
133 \DeclareInputMath{250}{\div}
134 \DeclareInputMath{251}{\pm}
135 \DeclareInputText{252}{\CYRNo}
136 \DeclareInputText{253}{\textcurrency}
137 \DeclareInputText{254}{\textblacksquare}

138 ⟨/AV⟩

The following block corresponds to the so called ‘Modified Alternative Variant’ (MAV) of cp 866. Symbols 0xf2 through 0xfd match standard IBM coding (MS code page 437):

139 ⟨∗MAV⟩

140 \DeclareInputMath{242}{\geq}
141 \DeclareInputMath{243}{\leq}
142 % 0xf4 TOP HALF INTEGRAL
143 % 0xf5 BOTTOM HALF INTEGRAL
144 \DeclareInputMath{246}{\div}
145 \DeclareInputMath{247}{\sim}
146 \DeclareInputMath{248}{\mathdegree}
147 \DeclareInputText{249}{\textbullet}
148 \DeclareInputText{250}{\textperiodcentered}
149 \DeclareInputMath{251}{\surd}
150 \DeclareInputMath{252}{\mathnsuperior}
151 \DeclareInputMath{253}{\mathtwosuperior}
152 \DeclareInputText{254}{\textblacksquare}

153 ⟨/MAV⟩

The following block corresponds to the yet another modern modification of cp 866:

154 ⟨∗NAV⟩

155 \DeclareInputText{242}{\CYRGUP}
156 \DeclareInputText{243}{\cyrgup}
157 \DeclareInputText{244}{\CYRIE}
158 \DeclareInputText{245}{\cyrie}
159 \DeclareInputText{246}{\CYRII}
160 \DeclareInputText{247}{\cyrii}
161 \DeclareInputText{248}{\CYRYI}
162 \DeclareInputText{249}{\cyryi}
163 \DeclareInputText{250}{\CYRUSHRT}
164 \DeclareInputText{251}{\cyrushrt}
165 \DeclareInputText{252}{\CYRNo}
166 % ? left European quotes:
167 \DeclareInputText{253}{\guillemotleft}
168 % ? right European quotes:
169 \DeclareInputText{254}{\guillemotright}

170 ⟨/NAV⟩ 171 %

172 \DeclareInputText{255}{\nobreakspace}

173 ⟨/cp866⟩

1.3 Microsoft cp 855

174 ⟨∗cp855⟩

175 \DeclareInputText{128}{\cyrdje}
176 \DeclareInputText{129}{\CYRDJE}
177 \DeclareInputText{130}{\@tabacckludge'\cyrg}
178 \DeclareInputText{131}{\@tabacckludge'\CYRG}
179 \DeclareInputText{132}{\cyryo}
180 \DeclareInputText{133}{\CYRYO}
181 \DeclareInputText{134}{\cyrie}
182 \DeclareInputText{135}{\CYRIE}
183 \DeclareInputText{136}{\cyrdze}

184 \DeclareInputText{137}{\CYRDZE}
185 \DeclareInputText{138}{\cyrii}
186 \DeclareInputText{139}{\CYRII}
187 \DeclareInputText{140}{\cyryi}
188 \DeclareInputText{141}{\CYRYI}
189 \DeclareInputText{142}{\cyrje}
190 \DeclareInputText{143}{\CYRJE}
191 \DeclareInputText{144}{\cyrlje}
192 \DeclareInputText{145}{\CYRLJE}
193 \DeclareInputText{146}{\cyrnje}
194 \DeclareInputText{147}{\CYRNJE}
195 \DeclareInputText{148}{\cyrtshe}
196 \DeclareInputText{149}{\CYRTSHE}
197 \DeclareInputText{150}{\@tabacckludge'\cyrk}
198 \DeclareInputText{151}{\@tabacckludge'\CYRK}
199 \DeclareInputText{152}{\cyrushrt}
200 \DeclareInputText{153}{\CYRUSHRT}
201 \DeclareInputText{154}{\cyrdzhe}
202 \DeclareInputText{155}{\CYRDZHE}
203 \DeclareInputText{156}{\cyryu}
204 \DeclareInputText{157}{\CYRYU}
205 \DeclareInputText{158}{\cyrhrdsn}
206 \DeclareInputText{159}{\CYRHRDSN}
207 \DeclareInputText{160}{\cyra}
208 \DeclareInputText{161}{\CYRA}
209 \DeclareInputText{162}{\cyrb}
210 \DeclareInputText{163}{\CYRB}
211 \DeclareInputText{164}{\cyrc}
212 \DeclareInputText{165}{\CYRC}
213 \DeclareInputText{166}{\cyrd}
214 \DeclareInputText{167}{\CYRD}
215 \DeclareInputText{168}{\cyre}
216 \DeclareInputText{169}{\CYRE}
217 \DeclareInputText{170}{\cyrf}
218 \DeclareInputText{171}{\CYRF}
219 \DeclareInputText{172}{\cyrg}
220 \DeclareInputText{173}{\CYRG}
221 \DeclareInputText{174}{\guillemotleft}
222 \DeclareInputText{175}{\guillemotright}
223 % 0xb0 LIGHT SHADE
224 % 0xb1 MEDIUM SHADE
225 % 0xb2 DARK SHADE
226 % 0xb3 BOX DRAWINGS LIGHT VERTICAL
227 % 0xb4 BOX DRAWINGS LIGHT VERTICAL AND LEFT
228 \DeclareInputText{181}{\cyrh}
229 \DeclareInputText{182}{\CYRH}
230 \DeclareInputText{183}{\cyri}
231 \DeclareInputText{184}{\CYRI}
232 % 0xb9 BOX DRAWINGS DOUBLE VERTICAL AND LEFT
233 % 0xba BOX DRAWINGS DOUBLE VERTICAL

234 % 0xbb BOX DRAWINGS DOUBLE DOWN AND LEFT
235 % 0xbc BOX DRAWINGS DOUBLE UP AND LEFT
236 \DeclareInputText{189}{\cyrishrt}
237 \DeclareInputText{190}{\CYRISHRT}
238 % 0xbf BOX DRAWINGS LIGHT DOWN AND LEFT
239 % 0xc0 BOX DRAWINGS LIGHT UP AND RIGHT
240 % 0xc1 BOX DRAWINGS LIGHT UP AND HORIZONTAL
241 % 0xc2 BOX DRAWINGS LIGHT DOWN AND HORIZONTAL
242 % 0xc3 BOX DRAWINGS LIGHT VERTICAL AND RIGHT
243 % 0xc4 BOX DRAWINGS LIGHT HORIZONTAL
244 % 0xc5 BOX DRAWINGS LIGHT VERTICAL AND HORIZONTAL
245 \DeclareInputText{198}{\cyrk}
246 \DeclareInputText{199}{\CYRK}
247 % 0xc8 BOX DRAWINGS DOUBLE UP AND RIGHT
248 % 0xc9 BOX DRAWINGS DOUBLE DOWN AND RIGHT
249 % 0xca BOX DRAWINGS DOUBLE UP AND HORIZONTAL
250 % 0xcb BOX DRAWINGS DOUBLE DOWN AND HORIZONTAL
251 % 0xcc BOX DRAWINGS DOUBLE VERTICAL AND RIGHT
252 % 0xcd BOX DRAWINGS DOUBLE HORIZONTAL
253 % 0xce BOX DRAWINGS DOUBLE VERTICAL AND HORIZONTAL
254 \DeclareInputText{207}{\textcurrency}
255 \DeclareInputText{208}{\cyrl}
256 \DeclareInputText{209}{\CYRL}
257 \DeclareInputText{210}{\cyrm}
258 \DeclareInputText{211}{\CYRM}
259 \DeclareInputText{212}{\cyrn}
260 \DeclareInputText{213}{\CYRN}
261 \DeclareInputText{214}{\cyro}
262 \DeclareInputText{215}{\CYRO}
263 \DeclareInputText{216}{\cyrp}
264 % 0xd9 BOX DRAWINGS LIGHT UP AND LEFT
265 % 0xda BOX DRAWINGS LIGHT DOWN AND RIGHT
266 % 0xdb FULL BLOCK
267 % 0xdc LOWER HALF BLOCK
268 \DeclareInputText{221}{\CYRP}
269 \DeclareInputText{222}{\cyrya}
270 % 0xdf UPPER HALF BLOCK
271 \DeclareInputText{224}{\CYRYA}
272 \DeclareInputText{225}{\cyrr}
273 \DeclareInputText{226}{\CYRR}
274 \DeclareInputText{227}{\cyrs}
275 \DeclareInputText{228}{\CYRS}
276 \DeclareInputText{229}{\cyrt}
277 \DeclareInputText{230}{\CYRT}
278 \DeclareInputText{231}{\cyru}
279 \DeclareInputText{232}{\CYRU}
280 \DeclareInputText{233}{\cyrzh}
281 \DeclareInputText{234}{\CYRZH}
282 \DeclareInputText{235}{\cyrv}
283 \DeclareInputText{236}{\CYRV}

284 \DeclareInputText{237}{\cyrsftsn}
285 \DeclareInputText{238}{\CYRSFTSN}
286 \DeclareInputText{239}{\CYRNo}
287 \DeclareInputText{240}{\-}
288 \DeclareInputText{241}{\cyrery}
289 \DeclareInputText{242}{\CYRERY}
290 \DeclareInputText{243}{\cyrz}
291 \DeclareInputText{244}{\CYRZ}
292 \DeclareInputText{245}{\cyrsh}
293 \DeclareInputText{246}{\CYRSH}
294 \DeclareInputText{247}{\cyrerev}
295 \DeclareInputText{248}{\CYREREV}
296 \DeclareInputText{249}{\cyrshch}
297 \DeclareInputText{250}{\CYRSHCH}
298 \DeclareInputText{251}{\cyrch}
299 \DeclareInputText{252}{\CYRCH}
300 \DeclareInputText{253}{\S}
301 \DeclareInputText{254}{\textblacksquare}
302 \DeclareInputText{255}{\nobreakspace}

303 ⟨/cp855⟩

1.4 Microsoft cp 1251

304 ⟨∗cp1251⟩

305 \DeclareInputText{192}{\CYRA}
306 \DeclareInputText{193}{\CYRB}
307 \DeclareInputText{194}{\CYRV}
308 \DeclareInputText{195}{\CYRG}
309 \DeclareInputText{196}{\CYRD}
310 \DeclareInputText{197}{\CYRE}
311 \DeclareInputText{198}{\CYRZH}
312 \DeclareInputText{199}{\CYRZ}
313 \DeclareInputText{200}{\CYRI}
314 \DeclareInputText{201}{\CYRISHRT}
315 \DeclareInputText{202}{\CYRK}
316 \DeclareInputText{203}{\CYRL}
317 \DeclareInputText{204}{\CYRM}
318 \DeclareInputText{205}{\CYRN}
319 \DeclareInputText{206}{\CYRO}
320 \DeclareInputText{207}{\CYRP}
321 \DeclareInputText{208}{\CYRR}
322 \DeclareInputText{209}{\CYRS}
323 \DeclareInputText{210}{\CYRT}
324 \DeclareInputText{211}{\CYRU}
325 \DeclareInputText{212}{\CYRF}
326 \DeclareInputText{213}{\CYRH}
327 \DeclareInputText{214}{\CYRC}
328 \DeclareInputText{215}{\CYRCH}
329 \DeclareInputText{216}{\CYRSH}
330 \DeclareInputText{217}{\CYRSHCH}

331 \DeclareInputText{218}{\CYRHRDSN}
332 \DeclareInputText{219}{\CYRERY}
333 \DeclareInputText{220}{\CYRSFTSN}
334 \DeclareInputText{221}{\CYREREV}
335 \DeclareInputText{222}{\CYRYU}
336 \DeclareInputText{223}{\CYRYA}

337 %

338 \DeclareInputText{224}{\cyra}
339 \DeclareInputText{225}{\cyrb}
340 \DeclareInputText{226}{\cyrv}
341 \DeclareInputText{227}{\cyrg}
342 \DeclareInputText{228}{\cyrd}
343 \DeclareInputText{229}{\cyre}
344 \DeclareInputText{230}{\cyrzh}
345 \DeclareInputText{231}{\cyrz}
346 \DeclareInputText{232}{\cyri}
347 \DeclareInputText{233}{\cyrishrt}
348 \DeclareInputText{234}{\cyrk}
349 \DeclareInputText{235}{\cyrl}
350 \DeclareInputText{236}{\cyrm}
351 \DeclareInputText{237}{\cyrn}
352 \DeclareInputText{238}{\cyro}
353 \DeclareInputText{239}{\cyrp}
354 \DeclareInputText{240}{\cyrr}
355 \DeclareInputText{241}{\cyrs}
356 \DeclareInputText{242}{\cyrt}
357 \DeclareInputText{243}{\cyru}
358 \DeclareInputText{244}{\cyrf}
359 \DeclareInputText{245}{\cyrh}
360 \DeclareInputText{246}{\cyrc}
361 \DeclareInputText{247}{\cyrch}
362 \DeclareInputText{248}{\cyrsh}
363 \DeclareInputText{249}{\cyrshch}
364 \DeclareInputText{250}{\cyrhrdsn}
365 \DeclareInputText{251}{\cyrery}
366 \DeclareInputText{252}{\cyrsftsn}
367 \DeclareInputText{253}{\cyrerev}
368 \DeclareInputText{254}{\cyryu}
369 \DeclareInputText{255}{\cyrya}

370 %

371 \DeclareInputText{128}{\CYRDJE}
372 \DeclareInputText{129}{\@tabacckludge'\CYRG}
373 \DeclareInputText{130}{\quotesinglbase}
374 \DeclareInputText{131}{\@tabacckludge'\cyrg}
375 \DeclareInputText{132}{\quotedblbase}
376 \DeclareInputText{133}{\dots}
377 \DeclareInputText{134}{\dag}
378 \DeclareInputText{135}{\ddag}
379 \DeclareInputText{136}{\texteuro}
380 \DeclareInputText{137}{\perthousand}

381 \DeclareInputText{138}{\CYRLJE}
382 \DeclareInputText{139}{\guilsinglleft}
383 \DeclareInputText{140}{\CYRNJE}
384 \DeclareInputText{141}{\@tabacckludge'\CYRK}
385 \DeclareInputText{142}{\CYRTSHE}
386 \DeclareInputText{143}{\CYRDZHE}
387 \DeclareInputText{144}{\cyrdje}
388 \DeclareInputText{145}{\textquoteleft}
389 \DeclareInputText{146}{\textquoteright}
390 \DeclareInputText{147}{\textquotedblleft}
391 \DeclareInputText{148}{\textquotedblright}
392 \DeclareInputText{149}{\textbullet}
393 \DeclareInputText{150}{\textendash}
394 \DeclareInputText{151}{\textemdash}
395 % 0x98 undefined
396 \DeclareInputText{153}{\texttrademark}
397 \DeclareInputText{154}{\cyrlje}
398 \DeclareInputText{155}{\guilsinglright}
399 \DeclareInputText{156}{\cyrnje}
400 \DeclareInputText{157}{\@tabacckludge'\cyrk}
401 \DeclareInputText{158}{\cyrtshe}
402 \DeclareInputText{159}{\cyrdzhe}
403 \DeclareInputText{160}{\nobreakspace}
404 \DeclareInputText{161}{\CYRUSHRT}
405 \DeclareInputText{162}{\cyrushrt}
406 \DeclareInputText{163}{\CYRJE}
407 \DeclareInputText{164}{\textcurrency}
408 \DeclareInputText{165}{\CYRGUP}
409 \DeclareInputText{166}{\textbrokenbar}
410 \DeclareInputText{167}{\S}
411 \DeclareInputText{168}{\CYRYO}
412 \DeclareInputText{169}{\textcopyright}
413 \DeclareInputText{170}{\CYRIE}
414 \DeclareInputText{171}{\guillemotleft}
415 \DeclareInputMath{172}{\lnot}
416 \DeclareInputText{173}{\-}
417 \DeclareInputText{174}{\textregistered}
418 \DeclareInputText{175}{\CYRYI}
419 \DeclareInputMath{176}{\mathdegree}
420 \DeclareInputMath{177}{\pm}
421 \DeclareInputText{178}{\CYRII}
422 \DeclareInputText{179}{\cyrii}
423 \DeclareInputText{180}{\cyrgup}
424 \DeclareInputMath{181}{\mu}
425 \DeclareInputText{182}{\P}
426 \DeclareInputText{183}{\textperiodcentered}
427 \DeclareInputText{184}{\cyryo}
428 \DeclareInputText{185}{\CYRNo}
429 \DeclareInputText{186}{\cyrie}
430 \DeclareInputText{187}{\guillemotright}

431 \DeclareInputText{188}{\cyrje}
432 \DeclareInputText{189}{\CYRDZE}
433 \DeclareInputText{190}{\cyrdze}
434 \DeclareInputText{191}{\cyryi}

435 ⟨/cp1251⟩

1.5 koi8

436 ⟨∗koi8⟩

437 \DeclareInputText{225}{\CYRA}
438 \DeclareInputText{226}{\CYRB}
439 \DeclareInputText{247}{\CYRV}
440 \DeclareInputText{231}{\CYRG}
441 \DeclareInputText{228}{\CYRD}
442 \DeclareInputText{229}{\CYRE}
443 \DeclareInputText{179}{\CYRYO}
444 \DeclareInputText{246}{\CYRZH}
445 \DeclareInputText{250}{\CYRZ}
446 \DeclareInputText{233}{\CYRI}
447 \DeclareInputText{234}{\CYRISHRT}
448 \DeclareInputText{235}{\CYRK}
449 \DeclareInputText{236}{\CYRL}
450 \DeclareInputText{237}{\CYRM}
451 \DeclareInputText{238}{\CYRN}
452 \DeclareInputText{239}{\CYRO}
453 \DeclareInputText{240}{\CYRP}
454 \DeclareInputText{242}{\CYRR}
455 \DeclareInputText{243}{\CYRS}
456 \DeclareInputText{244}{\CYRT}
457 \DeclareInputText{245}{\CYRU}
458 \DeclareInputText{230}{\CYRF}
459 \DeclareInputText{232}{\CYRH}
460 \DeclareInputText{227}{\CYRC}
461 \DeclareInputText{254}{\CYRCH}
462 \DeclareInputText{251}{\CYRSH}
463 \DeclareInputText{253}{\CYRSHCH}
464 \DeclareInputText{255}{\CYRHRDSN}
465 \DeclareInputText{249}{\CYRERY}
466 \DeclareInputText{248}{\CYRSFTSN}
467 \DeclareInputText{252}{\CYREREV}
468 \DeclareInputText{224}{\CYRYU}
469 \DeclareInputText{241}{\CYRYA}

470 %

471 \DeclareInputText{193}{\cyra}
472 \DeclareInputText{194}{\cyrb}
473 \DeclareInputText{215}{\cyrv}
474 \DeclareInputText{199}{\cyrg}
475 \DeclareInputText{196}{\cyrd}
476 \DeclareInputText{197}{\cyre}
477 \DeclareInputText{163}{\cyryo}

478 \DeclareInputText{214}{\cyrzh}
479 \DeclareInputText{218}{\cyrz}
480 \DeclareInputText{201}{\cyri}
481 \DeclareInputText{202}{\cyrishrt}
482 \DeclareInputText{203}{\cyrk}
483 \DeclareInputText{204}{\cyrl}
484 \DeclareInputText{205}{\cyrm}
485 \DeclareInputText{206}{\cyrn}
486 \DeclareInputText{207}{\cyro}
487 \DeclareInputText{208}{\cyrp}
488 \DeclareInputText{210}{\cyrr}
489 \DeclareInputText{211}{\cyrs}
490 \DeclareInputText{212}{\cyrt}
491 \DeclareInputText{213}{\cyru}
492 \DeclareInputText{198}{\cyrf}
493 \DeclareInputText{200}{\cyrh}
494 \DeclareInputText{195}{\cyrc}
495 \DeclareInputText{222}{\cyrch}
496 \DeclareInputText{219}{\cyrsh}
497 \DeclareInputText{221}{\cyrshch}
498 \DeclareInputText{223}{\cyrhrdsn}
499 \DeclareInputText{217}{\cyrery}
500 \DeclareInputText{216}{\cyrsftsn}
501 \DeclareInputText{220}{\cyrerev}
502 \DeclareInputText{192}{\cyryu}
503 \DeclareInputText{209}{\cyrya}

1.5.1 koi8-r and relatives (koi8-ru, koi8-u, ISO-IR-111)

0x80–0x9f are unused in ISO-IR-111 ECMA Cyrillic Code Page

504 ⟨∗koi8r | koi8ru⟩

505 % 0x80 FORMS LIGHT HORIZONTAL
506 % 0x81 FORMS LIGHT VERTICAL
507 % 0x82 FORMS LIGHT DOWN AND RIGHT
508 % 0x83 FORMS LIGHT DOWN AND LEFT
509 % 0x84 FORMS LIGHT UP AND RIGHT
510 % 0x85 FORMS LIGHT UP AND LEFT
511 % 0x86 FORMS LIGHT VERTICAL AND RIGHT
512 % 0x87 FORMS LIGHT VERTICAL AND LEFT
513 % 0x88 FORMS LIGHT DOWN AND HORIZONTAL
514 % 0x89 FORMS LIGHT UP AND HORIZONTAL
515 % 0x8A FORMS LIGHT VERTICAL AND HORIZONTAL
516 % 0x8B UPPER HALF BLOCK
517 % 0x8C LOWER HALF BLOCK
518 % 0x8D FULL BLOCK
519 % 0x8E LEFT HALF BLOCK
520 % 0x8F RIGHT HALF BLOCK
521 % 0x90 LIGHT SHADE
522 % 0x91 MEDIUM SHADE
523 % 0x92 DARK SHADE

524 ⟨/koi8r | koi8ru⟩ 525 ⟨∗koi8r⟩

526 % 0x93 TOP HALF INTEGRAL

527 ⟨/koi8r⟩

528 ⟨koi8ru⟩\DeclareInputText{147}{\quotedblleft}

529 ⟨∗koi8r | koi8ru⟩

530 \DeclareInputText{148}{\textblacksquare}
531 \DeclareInputText{149}{\textbullet}

532 ⟨/koi8r | koi8ru⟩ 533 ⟨∗koi8r⟩

534 \DeclareInputMath{150}{\surd}
535 \DeclareInputMath{151}{\sim}
536 \DeclareInputMath{152}{\leq}
537 \DeclareInputMath{153}{\geq}

538 ⟨/koi8r⟩ 539 ⟨∗koi8ru⟩

540 \DeclareInputText{150}{\quotedblright}
541 \DeclareInputText{151}{\textemdash}
542 \DeclareInputText{152}{\CYRNo}
543 \DeclareInputText{153}{\texttrademark}

544 ⟨/koi8ru⟩

545 ⟨koi8r | koi8ru⟩\DeclareInputText{154}{\nobreakspace}

546 ⟨∗koi8r⟩

547 % 0x9B BOTTOM HALF INTEGRAL
548 \DeclareInputMath{156}{\mathdegree}
549 \DeclareInputMath{157}{\mathtwosuperior}

550 ⟨/koi8r⟩ 551 ⟨∗koi8ru⟩

552 \DeclareInputText{155}{\guillemotright}
553 \DeclareInputText{156}{\textregistered}
554 \DeclareInputText{157}{\guillemotleft}

555 ⟨/koi8ru⟩

556 ⟨koi8r | koi8ru⟩\DeclareInputText{158}{\textperiodcentered}
557 ⟨koi8r⟩\DeclareInputMath{159}{\div}
558 ⟨koi8ru⟩\DeclareInputText{159}{\textcurrency}

559 ⟨∗koi8r | koi8ru⟩

560 % 0xA0 FORMS DOUBLE HORIZONTAL
561 % 0xA1 FORMS DOUBLE VERTICAL
562 % 0xA2 FORMS DOWN SINGLE AND RIGHT DOUBLE

563 ⟨/koi8r | koi8ru⟩ 564 ⟨∗isoir111⟩

565 \DeclareInputText{160}{\nobreakspace}
566 \DeclareInputText{161}{\cyrdje}
567 \DeclareInputText{162}{\@tabacckludge'\cyrg}

568 ⟨/isoir111⟩ 569 ⟨∗koi8r⟩

570 % 0xA4 FORMS DOWN DOUBLE AND RIGHT SINGLE

571 ⟨/koi8r⟩

572 ⟨koi8ru | isoir111 | koi8u⟩\DeclareInputText{164}{\cyrie}

573 ⟨∗koi8r | koi8ru⟩

574 % 0xA5 FORMS DOUBLE DOWN AND RIGHT

575 ⟨/koi8r | koi8ru⟩

576 ⟨isoir111⟩\DeclareInputText{165}{\cyrdze}

577 ⟨∗koi8r⟩

578 % 0xA6 FORMS DOWN SINGLE AND LEFT DOUBLE
579 % 0xA7 FORMS DOWN DOUBLE AND LEFT SINGLE

580 ⟨/koi8r⟩ 581 ⟨∗koi8ru | isoir111 | koi8u⟩

582 \DeclareInputText{166}{\cyrii}
583 \DeclareInputText{167}{\cyryi}

584 ⟨/koi8ru | isoir111 | koi8u⟩ 585 ⟨∗koi8r | koi8ru⟩

586 % 0xA8 FORMS DOUBLE DOWN AND LEFT
587 % 0xA9 FORMS UP SINGLE AND RIGHT DOUBLE
588 % 0xAA FORMS UP DOUBLE AND RIGHT SINGLE
589 % 0xAB FORMS DOUBLE UP AND RIGHT
590 % 0xAC FORMS UP SINGLE AND LEFT DOUBLE

591 ⟨/koi8r | koi8ru⟩ 592 ⟨∗isoir111⟩

593 \DeclareInputText{168}{\cyrje}
594 \DeclareInputText{169}{\cyrlje}
595 \DeclareInputText{170}{\cyrnje}
596 \DeclareInputText{171}{\cyrtshe}
597 \DeclareInputText{172}{\@tabacckludge'\cyrk}

598 ⟨/isoir111⟩ 599 ⟨∗koi8r⟩

600 % 0xAD FORMS UP DOUBLE AND LEFT SINGLE
601 % 0xAE FORMS DOUBLE UP AND LEFT

602 ⟨/koi8r⟩

603 ⟨koi8ru | koi8u⟩\DeclareInputText{173}{\cyrgup}
604 ⟨isoir111⟩\DeclareInputText{173}{\-}
605 ⟨koi8ru | isoir111⟩\DeclareInputText{174}{\cyrushrt}

606 ⟨∗koi8r | koi8ru⟩

607 % 0xAF FORMS VERTICAL SINGLE AND RIGHT DOUBLE
608 % 0xB0 FORMS VERTICAL DOUBLE AND RIGHT SINGLE
609 % 0xB1 FORMS DOUBLE VERTICAL AND RIGHT
610 % 0xB2 FORMS VERTICAL SINGLE AND LEFT DOUBLE

611 ⟨/koi8r | koi8ru⟩ 612 ⟨∗isoir111⟩

613 \DeclareInputText{175}{\cyrdzhe}
614 \DeclareInputText{176}{\CYRNo}
615 \DeclareInputText{177}{\CYRDJE}
616 \DeclareInputText{178}{\@tabacckludge'CYRG}

617 ⟨/isoir111⟩ 618 ⟨∗koi8r⟩

619 % 0xB4 FORMS VERTICAL DOUBLE AND LEFT SINGLE

620 ⟨/koi8r⟩

621 ⟨koi8ru | isoir111 | koi8u⟩\DeclareInputText{180}{\CYRIE}

622 ⟨∗koi8r | koi8ru⟩

623 % 0xB5 FORMS DOUBLE VERTICAL AND LEFT

624 ⟨/koi8r | koi8ru⟩

625 ⟨isoir111⟩\DeclareInputText{181}{\CYRDZE}

626 ⟨∗koi8r⟩

627 % 0xB6 FORMS DOWN SINGLE AND HORIZONTAL DOUBLE
628 % 0xB7 FORMS DOWN DOUBLE AND HORIZONTAL SINGLE

629 ⟨/koi8r⟩ 630 ⟨∗koi8ru | isoir111 | koi8u⟩

631 \DeclareInputText{182}{\CYRII}
632 \DeclareInputText{183}{\CYRYI}

633 ⟨/koi8ru | isoir111 | koi8u⟩ 634 ⟨∗koi8r | koi8ru⟩

635 % 0xB8 FORMS DOUBLE DOWN AND HORIZONTAL
636 % 0xB9 FORMS UP SINGLE AND HORIZONTAL DOUBLE
637 % 0xBA FORMS UP DOUBLE AND HORIZONTAL SINGLE
638 % 0xBB FORMS DOUBLE UP AND HORIZONTAL
639 % 0xBC FORMS VERTICAL SINGLE AND HORIZONTAL DOUBLE

640 ⟨/koi8r | koi8ru⟩ 641 ⟨∗isoir111⟩

642 \DeclareInputText{184}{\CYRJE}
643 \DeclareInputText{185}{\CYRLJE}
644 \DeclareInputText{186}{\CYRNJE}
645 \DeclareInputText{187}{\CYRTSHE}
646 \DeclareInputText{188}{\@tabacckludge'\CYRK}

647 ⟨/isoir111⟩ 648 ⟨∗koi8r⟩

649 % 0xBD FORMS VERTICAL DOUBLE AND HORIZONTAL SINGLE
650 % 0xBE FORMS DOUBLE VERTICAL AND HORIZONTAL

651 ⟨/koi8r⟩

652 ⟨koi8ru | koi8u⟩\DeclareInputText{189}{\CYRGUP}
653 ⟨isoir111⟩\DeclareInputText{189}{\textcurrency}
654 ⟨koi8ru | isoir111⟩\DeclareInputText{190}{\CYRUSHRT}
655 ⟨koi8r | koi8ru⟩\DeclareInputText{191}{\textcopyright}
656 ⟨isoir111⟩\DeclareInputText{191}{\CYRDZHE}

657 ⟨/koi8⟩

1.6 ISO 8859-5

658 ⟨∗ISO88595⟩

659 \DeclareInputText{160}{\nobreakspace}
660 \DeclareInputText{161}{\CYRYO}
661 \DeclareInputText{162}{\CYRDJE}
662 \DeclareInputText{163}{\@tabacckludge'\CYRG}
663 \DeclareInputText{164}{\CYRIE}
664 \DeclareInputText{165}{\CYRDZE}
665 \DeclareInputText{166}{\CYRII}
666 \DeclareInputText{167}{\CYRYI}
667 \DeclareInputText{168}{\CYRJE}
668 \DeclareInputText{169}{\CYRLJE}
669 \DeclareInputText{170}{\CYRNJE}
670 \DeclareInputText{171}{\CYRTSHE}

671 \DeclareInputText{172}{\@tabacckludge'\CYRK}
672 \DeclareInputText{173}{\-}
673 \DeclareInputText{174}{\CYRUSHRT}
674 \DeclareInputText{175}{\CYRDZHE}

675 %

676 \DeclareInputText{176}{\CYRA}
677 \DeclareInputText{177}{\CYRB}
678 \DeclareInputText{178}{\CYRV}
679 \DeclareInputText{179}{\CYRG}
680 \DeclareInputText{180}{\CYRD}
681 \DeclareInputText{181}{\CYRE}
682 \DeclareInputText{182}{\CYRZH}
683 \DeclareInputText{183}{\CYRZ}
684 \DeclareInputText{184}{\CYRI}
685 \DeclareInputText{185}{\CYRISHRT}
686 \DeclareInputText{186}{\CYRK}
687 \DeclareInputText{187}{\CYRL}
688 \DeclareInputText{188}{\CYRM}
689 \DeclareInputText{189}{\CYRN}
690 \DeclareInputText{190}{\CYRO}
691 \DeclareInputText{191}{\CYRP}
692 \DeclareInputText{192}{\CYRR}
693 \DeclareInputText{193}{\CYRS}
694 \DeclareInputText{194}{\CYRT}
695 \DeclareInputText{195}{\CYRU}
696 \DeclareInputText{196}{\CYRF}
697 \DeclareInputText{197}{\CYRH}
698 \DeclareInputText{198}{\CYRC}
699 \DeclareInputText{199}{\CYRCH}
700 \DeclareInputText{200}{\CYRSH}
701 \DeclareInputText{201}{\CYRSHCH}
702 \DeclareInputText{202}{\CYRHRDSN}
703 \DeclareInputText{203}{\CYRERY}
704 \DeclareInputText{204}{\CYRSFTSN}
705 \DeclareInputText{205}{\CYREREV}
706 \DeclareInputText{206}{\CYRYU}
707 \DeclareInputText{207}{\CYRYA}

708 %

709 \DeclareInputText{208}{\cyra}
710 \DeclareInputText{209}{\cyrb}
711 \DeclareInputText{210}{\cyrv}
712 \DeclareInputText{211}{\cyrg}
713 \DeclareInputText{212}{\cyrd}
714 \DeclareInputText{213}{\cyre}
715 \DeclareInputText{214}{\cyrzh}
716 \DeclareInputText{215}{\cyrz}
717 \DeclareInputText{216}{\cyri}
718 \DeclareInputText{217}{\cyrishrt}
719 \DeclareInputText{218}{\cyrk}
720 \DeclareInputText{219}{\cyrl}

721 \DeclareInputText{220}{\cyrm}
722 \DeclareInputText{221}{\cyrn}
723 \DeclareInputText{222}{\cyro}
724 \DeclareInputText{223}{\cyrp}
725 \DeclareInputText{224}{\cyrr}
726 \DeclareInputText{225}{\cyrs}
727 \DeclareInputText{226}{\cyrt}
728 \DeclareInputText{227}{\cyru}
729 \DeclareInputText{228}{\cyrf}
730 \DeclareInputText{229}{\cyrh}
731 \DeclareInputText{230}{\cyrc}
732 \DeclareInputText{231}{\cyrch}
733 \DeclareInputText{232}{\cyrsh}
734 \DeclareInputText{233}{\cyrshch}
735 \DeclareInputText{234}{\cyrhrdsn}
736 \DeclareInputText{235}{\cyrery}
737 \DeclareInputText{236}{\cyrsftsn}
738 \DeclareInputText{237}{\cyrerev}
739 \DeclareInputText{238}{\cyryu}
740 \DeclareInputText{239}{\cyrya}

741 %

742 \DeclareInputText{240}{\CYRNo}
743 \DeclareInputText{241}{\cyryo}
744 \DeclareInputText{242}{\cyrdje}
745 \DeclareInputText{243}{\@tabacckludge'\cyrg}
746 \DeclareInputText{244}{\cyrie}
747 \DeclareInputText{245}{\cyrdze}
748 \DeclareInputText{246}{\cyrii}
749 \DeclareInputText{247}{\cyryi}
750 \DeclareInputText{248}{\cyrje}
751 \DeclareInputText{249}{\cyrlje}
752 \DeclareInputText{250}{\cyrnje}
753 \DeclareInputText{251}{\cyrtshe}
754 \DeclareInputText{252}{\@tabacckludge'\cyrk}
755 \DeclareInputText{253}{\S}
756 \DeclareInputText{254}{\cyrushrt}
757 \DeclareInputText{255}{\cyrdzhe}

758 ⟨/ISO88595⟩

1.7 Cyrillic Mac encodings

The MacOS Cyrillic encoding (Microsoft cp 10007) includes the full Cyrillic letter repertoire of ISO 8859-5 (although not at the same code points). This covers most of the Slavic languages written in Cyrillic script.

The MacOS Cyrillic encoding also includes a number of characters needed for the MacOS user interface (e.g., ellipsis, bullet for echoing passwords, copyright sign, etc). All of the characters in MacOS Cyrillic that are also in the MacOS Roman encoding are at the same code points as MacOS Roman. This improves application compatibility (since some naughty applications hard-code the MacOS Roman code points of certain characters).

A variant of MacOS Cyrillic is used for Ukrainian. This character set adds upper and lower GHE WITH UPTURN, for a grand total of 2 code point differences from standard MacOS Cyrillic.

759 ⟨∗maccyrillic | macukrainian⟩

760 \DeclareInputText{128}{\CYRA}
761 \DeclareInputText{129}{\CYRB}
762 \DeclareInputText{130}{\CYRV}
763 \DeclareInputText{131}{\CYRG}
764 \DeclareInputText{132}{\CYRD}
765 \DeclareInputText{133}{\CYRE}
766 \DeclareInputText{134}{\CYRZH}
767 \DeclareInputText{135}{\CYRZ}
768 \DeclareInputText{136}{\CYRI}
769 \DeclareInputText{137}{\CYRISHRT}
770 \DeclareInputText{138}{\CYRK}
771 \DeclareInputText{139}{\CYRL}
772 \DeclareInputText{140}{\CYRM}
773 \DeclareInputText{141}{\CYRN}
774 \DeclareInputText{142}{\CYRO}
775 \DeclareInputText{143}{\CYRP}
776 \DeclareInputText{144}{\CYRR}
777 \DeclareInputText{145}{\CYRS}
778 \DeclareInputText{146}{\CYRT}
779 \DeclareInputText{147}{\CYRU}
780 \DeclareInputText{148}{\CYRF}
781 \DeclareInputText{149}{\CYRH}
782 \DeclareInputText{150}{\CYRC}
783 \DeclareInputText{151}{\CYRCH}
784 \DeclareInputText{152}{\CYRSH}
785 \DeclareInputText{153}{\CYRSHCH}
786 \DeclareInputText{154}{\CYRHRDSN}
787 \DeclareInputText{155}{\CYRERY}
788 \DeclareInputText{156}{\CYRSFTSN}
789 \DeclareInputText{157}{\CYREREV}
790 \DeclareInputText{158}{\CYRYU}
791 \DeclareInputText{159}{\CYRYA}

792 %

793 \DeclareInputText{160}{\dag}
794 \DeclareInputMath{161}{\mathdegree}
795 ⟨maccyrillic⟩\DeclareInputText{162}{\textcent}
796 ⟨macukrainian⟩\DeclareInputText{162}{\CYRGUP}
797 \DeclareInputText{163}{\pounds}
798 \DeclareInputText{164}{\S}
799 \DeclareInputText{165}{\textbullet}
800 \DeclareInputText{166}{\P}
801 \DeclareInputText{167}{\CYRII}
802 \DeclareInputText{168}{\textregistered}
803 \DeclareInputText{169}{\textcopyright}
804 \DeclareInputText{170}{\texttrademark}

805 \DeclareInputText{171}{\CYRDJE}
806 \DeclareInputText{172}{\cyrdje}
807 \DeclareInputMath{173}{\neq}
808 \DeclareInputText{174}{\@tabacckludge'\CYRG}
809 \DeclareInputText{175}{\@tabacckludge'\cyrg}
810 \DeclareInputMath{176}{\infty}
811 \DeclareInputMath{177}{\pm}
812 \DeclareInputMath{178}{\leq}
813 \DeclareInputMath{179}{\geq}
814 \DeclareInputText{180}{\cyrii}
815 \DeclareInputMath{181}{\mu}
816 ⟨maccyrillic⟩\DeclareInputMath{182}{\partial}
817 ⟨macukrainian⟩\DeclareInputText{182}{\cyrgup}

818 %

819 \DeclareInputText{183}{\CYRJE}
820 \DeclareInputText{184}{\CYRIE}
821 \DeclareInputText{185}{\cyrie}
822 \DeclareInputText{186}{\CYRYI}
823 \DeclareInputText{187}{\cyryi}
824 \DeclareInputText{188}{\CYRLJE}
825 \DeclareInputText{189}{\cyrlje}
826 \DeclareInputText{190}{\CYRNJE}
827 \DeclareInputText{191}{\cyrnje}
828 \DeclareInputText{192}{\cyrje}
829 \DeclareInputText{193}{\CYRDZE}

830 %

831 \DeclareInputMath{194}{\lnot}
832 \DeclareInputMath{195}{\surd}
833 \DeclareInputText{196}{\textflorin}
834 \DeclareInputMath{197}{\approx}
835 % INCREMENT:
836 \DeclareInputMath{198}{\Delta}
837 \DeclareInputText{199}{\guillemotleft}
838 \DeclareInputText{200}{\guillemotright}
839 % HORIZONTAL ELLIPSIS:
840 \DeclareInputText{201}{\dots}
841 \DeclareInputText{202}{\nobreakspace}

842 %

843 \DeclareInputText{203}{\CYRTSHE}
844 \DeclareInputText{204}{\cyrtshe}
845 \DeclareInputText{205}{\@tabacckludge'\CYRK}
846 \DeclareInputText{206}{\@tabacckludge'\cyrk}
847 \DeclareInputText{207}{\cyrdze}

848 %

849 \DeclareInputText{208}{\textendash}
850 \DeclareInputText{209}{\textemdash}
851 \DeclareInputText{210}{\textquotedblleft}
852 \DeclareInputText{211}{\textquotedblright}
853 \DeclareInputText{212}{\textquoteleft}
854 \DeclareInputText{213}{\textquoteright}

855 \DeclareInputMath{214}{\div}
856 \DeclareInputText{215}{\quotedblbase}

857 %

858 \DeclareInputText{216}{\CYRUSHRT}
859 \DeclareInputText{217}{\cyrushrt}
860 \DeclareInputText{218}{\CYRDZHE}
861 \DeclareInputText{219}{\cyrdzhe}
862 \DeclareInputText{220}{\CYRNo}

863 %

864 \DeclareInputText{221}{\CYRYO}
865 \DeclareInputText{222}{\cyryo}
866 \DeclareInputText{223}{\cyrya}
867 \DeclareInputText{224}{\cyra}
868 \DeclareInputText{225}{\cyrb}
869 \DeclareInputText{226}{\cyrv}
870 \DeclareInputText{227}{\cyrg}
871 \DeclareInputText{228}{\cyrd}
872 \DeclareInputText{229}{\cyre}
873 \DeclareInputText{230}{\cyrzh}
874 \DeclareInputText{231}{\cyrz}
875 \DeclareInputText{232}{\cyri}
876 \DeclareInputText{233}{\cyrishrt}
877 \DeclareInputText{234}{\cyrk}
878 \DeclareInputText{235}{\cyrl}
879 \DeclareInputText{236}{\cyrm}
880 \DeclareInputText{237}{\cyrn}
881 \DeclareInputText{238}{\cyro}
882 \DeclareInputText{239}{\cyrp}
883 \DeclareInputText{240}{\cyrr}
884 \DeclareInputText{241}{\cyrs}
885 \DeclareInputText{242}{\cyrt}
886 \DeclareInputText{243}{\cyru}
887 \DeclareInputText{244}{\cyrf}
888 \DeclareInputText{245}{\cyrh}
889 \DeclareInputText{246}{\cyrc}
890 \DeclareInputText{247}{\cyrch}
891 \DeclareInputText{248}{\cyrsh}
892 \DeclareInputText{249}{\cyrshch}
893 \DeclareInputText{250}{\cyrhrdsn}
894 \DeclareInputText{251}{\cyrery}
895 \DeclareInputText{252}{\cyrsftsn}
896 \DeclareInputText{253}{\cyrerev}
897 \DeclareInputText{254}{\cyryu}
898 \DeclareInputText{255}{\textcurrency}

899 ⟨/maccyrillic | macukrainian⟩